### v1

# Daria Semochkina School of Mathematical Sciences

Abstract
Background Statisticians evaluating the impact of policy interventions such as screening or vaccination will need to make use of mathematical and computational models of disease progression and spread. Calibration is the process of identifying the parameters of these models, with a Bayesian framework providing a natural way in which to do this in a probabilistic fashion. Markov Chain Monte Carlo (MCMC) is one of a number of computational tools that is useful in carrying out this calibration. Objective In the context of complex models in particular, a key problem that arises is one of non-identifiability. In this setting, one approach which can be used is to consider and ensure that appropriately informative priors are specified on the joint parameter space. We give examples of how this arises and may be addressed in practice. Methods Using a basic SIS model the calibration process and the associated challenge of non-identifiability is discussed. How this problem arises in the context of a larger model for HPV and cervical cancer is also illustrated. Results The conditions which allow the problem of non-identifiability to be resolved are demonstrated for the SIS model. For the larger HPV model, how this impacts on the calibration process is also discussed.
Keywords MCMC, Metropolis-Hastings, Bayesian, model calibration, uncertainty quantification, human papillomavirus, HPV, cervical cancer, Ireland
1 Introduction
As statisticians, we are often involved as part of a larger team of applied mathematicians, computer modellers and epidemiologists when fitting and using disease models such as those described in this article.
The use of mathematical models to predict likely outcomes of different medical interventions, such as vaccination, screening programs and treatments to tackle a disease can be extremely helpful. Computer-based models of disease spread and progression are used for medical decision making, economic and policy evaluation in disease evaluation and control. In the case of health systems, decision makers are often doctors or administrators, making decisions on behalf of patients, current and future users and of the health system.
The key issue examined in this paper where the input of the statistician is important is the area of model calibration. We examine an approach to calibration, how the stochastic samples from the parameter sets are summarised and how to deal with a key challenging aspect, namely that of non-identifiability.
In healthcare applications, model calibration is the process of identifying a set or sets of parameters so that the mathematical model yields predictions for population characteristics (e.g. mortality, number of cases of a disease, the severity of illness) that match what is actually observed. As there is a great demand for model calibration techniques, there exist numerous methods and approaches [1].
In model calibration, so-called ‘calibration targets’ are usually chosen to correspond to some data that are observed directly. The aim of Bayesian model calibration is to find a distribution for unobservable parameters (model inputs), conditional on available data (calibration targets). This is often referred to as an inverse problem. In some situations, it may be feasible and sufficient to obtain a set of parameters using a best-fit approach or a pure optimisation technique. However, these approaches underestimate the uncertainty associated with any forecasts that may be made with the model by only providing point estimates. It is therefore desirable that this uncertainty in the distribution of the parameters obtained through calibration be quantified somehow. In addition, some frequently adopted approaches such as grid search produce worse fits of the model to the observed data when using the same computational resources[2].
The Bayesian framework is a natural one to use for solving inverse problems in general and consequently perfect for model calibration. The advantages of the Bayesian approach include conditional sampling and uncertainty quantification. The main computational complexity in Bayesian inversion is in sampling from a posterior distribution. The distribution of interest is often high-dimensional and does not have a standard form. Markov chain Monte Carlo (MCMC) methods are powerful tools for solving inverse problems and obtaining a sample from the posterior distribution of interest[2].
We focus this paper on solving inverse problems[3], illustrating it with a human papillomavirus (HPV) model calibration example. While other approaches exist for sampling from Bayesian posterior distributions, here we use MCMC algorithms, in particular the Metropolis-Hastings algorithm. This approach and this example is useful from the point of view of illustrating generic calibration problems, as well as being a realistic model for a topic of substantive interest.
The Bayesian approach involves obtaining a probability distribution for the parameters of interest, $\theta$, having seen some data, $\mathscr{D}$. In this case, the data are the ‘target’ observations, e.g. prevalence of disease at a particular time point. It is a necessary feature of the paradigm that ‘prior’ distributions for the parameters of interest are formulated, which encapsulate our understanding of the parameter space without directly incorporating evidence from the data.
Once a prior has been specified, and targets have been obtained, Bayes Theorem allows us to obtain an expression for the posterior distribution.
$\mathscr{P}(\theta|\mathscr{D})\propto\mathscr{P}(\mathscr{D}|\theta)\cdot\mathscr{P}(\theta),$
where $\mathscr{P(\theta)}$ is the prior distribution and $\mathscr{P}(\mathscr{D}|\theta)$ is the likelihood function. All of the information relevant to $\theta$ contained in the targets (here called $\mathscr{D}$) is contained in the likelihood function. Other information about our parameters can be summarised in the prior. Where there is no additional information the prior is often specified as a uniform distribution, in which case the posterior has the same shape as the likelihood. Of particular importance here is that information which is not contained in the targets, that is relevant to the inference may be incorporated through the prior. As will be seen later, a reconsideration about what external information about $\theta$ exists may be useful in addressing the question of non-identifiability.
In the next section, we discuss the methodology of calibration in general, introduce the Metropolis-Hastings algorithm and discuss potential limitations. Some approaches to resolve non-identifiability is discussed in Section 3. We then proceed to examine this in the context of a simple SIS disease transmission model in Section 4. We illustrate the role of the prior here in inducing identifiability during the calibration process. In Section 5, we present the HPV model and the Ireland-specific data to which the model is calibrated. In Section 6, we present simulation results with a few modifications of the Metropolis-Hastings algorithm to improve the performance, including an examination of the effects of having an informative prior distribution. Finally, Section 7 provides some discussion and conclusions.
2 Model calibration using the Metropolis-Hastings algorithm
If we formulate the model calibration problem in a probabilistic way, the output of model calibration will be a sample from the parameters’ a posteriori probability distribution. Assuming that each parameter combination produces a unique forward model causes the outcome of model calibration to not be a unique model but rather a collection of models. A single model could be the result of techniques that produce a single best-fitting point estimate in the parameter space. These models can be equally consistent with the observed data. This collection should agree with both the a priori information—incorporated in the model through prior distributions—and the observed data, accurately incorporating the uncertainty about observations. The Metropolis-Hastings algorithm searches through the joint parameter space until the chain converges in distribution around sets of values (the stationary distribution of the Markov chain is reached), which may then be used for inference or decision making.
The Metropolis-Hastings algorithm (M-H) is an MCMC algorithm for obtaining a random sample from a probability distribution $f$, for which direct sampling is difficult. The M-H algorithm has been tremendously successful in modern computational statistics because it enables sampling from complex distributions to obtain independent samples directly from such distributions. The algorithm is based on proposing values, sampled from some proposal distribution. These values are then accepted or rejected with a certain probability of acceptance, based on how likely those values are for some distribution of interest $f$. See Appendix A.1 for details.
2.1 Challenges
A number of challenges exist when calibrating models using limited data. These include difficulties related to the high dimensionality of the parameter space, multi-modality of the likelihood function and the question of identifiability.
When a very large number of parameters are present (i.e. high dimensionality) this increases the complexity and the amount of computation required regardless of the numerical algorithm used.
Where multiple modes (local maxima) are present in the likelihood, algorithms risk ‘getting stuck’ at local optima.
The challenge of identifiability[4] arises where the data are consistent with at least two (often a continuum of) different sets of parameter values. For example in the case of disease progression, the data could be consistent with situations where the rate of progression is fast and the rate of recovery is fast but equally, it could be the situation where the rate of recovery is slow with a slow rate of progression.
Some models (e.g. micro-simulation or agent based models) are computationally intensive, and so it is necessary to ensure the sampling algorithm is efficient
3 Approach to resolving non-identifiability
In practice, non-identifiability can be diagnosed when there is a lack of convergence of chains, when the posterior distributions of parameters are flat and broad, and when there is cross-correlation between sampled values.
In order to resolve non-identifiability, the suggested approach is to include information about some of the variables where this is obtainable. This can be done by carefully considering whether additional information has been collected or is available from other sources. For example duration of illness may not be available locally, but a prior may be obtained by considering international evidence that is emerging. In algorithmic terms the process of calibration would proceed as follows:
  • Identify targets and populate these with collected data.
  • Specify priors for parameters with a ‘first pass’, perhaps reasonably vague specification.
  • Run MCMC calibration process.
  • Identify parameters that are unidentifiable, together with those that are correlated.
  • Consider whether additional evidence may be obtained from external sources.
  • Quantify information about this in terms of a ‘prior’ probability distribution.
  • Rerun MCMC sampling to calibrate the model conditional on this additional evidence.
4 Example: SIS model
To illustrate this further we provide an example in line with the model introduced in Figure 1(a). We consider a simple susceptible-infected-susceptible (SIS) compartmental model. The transitions between the two states are described by
$\mathop{}\!\mathrm{d}S$ $=-(\beta\cdot S\cdot I)+\gamma\cdot I$
$\mathop{}\!\mathrm{d}I$ $=(\beta\cdot S\cdot I)-\gamma\cdot I,$
where $\beta=c\cdot p$ and $\gamma=1/d$ with $c$ being the contact rate, $p$ the transition probability and $d$ the infectious period. Most of the population starts off in the susceptible state with $S=0.99$ and $I=0.01$. We treat this situation as a population where there has initially been disease introduced, and at some later stage, a stable situation has been observed where 1200 of 2000 individuals are identified as infected. Thus, the likelihood may be described as having a similar shape to normal distributions $N(0.4,0.01)$ and $N(0.6,0.01)$, and these targets are then used in the algorithm. Implicit in this is that an improper uniform prior is used for p and c.
After running a simple Metropolis-Hastings algorithm for $c,\ p$ and $d$ with those normal distributions as our targets we plot the resulting $\beta$ and $\gamma$ values in Figure 1(b). It is noticeable that those are almost perfectly correlated. In panel (b), of the figure, the impact of specifying a prior for one variable is shown. Even though no information about the second variable is added, conditioning on a narrower prior distribution for $\gamma$ restricts the joint posterior space to a narrower range, thus impacting the marginal posterior distribution for $\beta$ which is induced due to the correlation between the parameters. Additionally, because $\beta=c\cdot p$, $c$ and $p$ are perfectly correlated. This leads to the chains for $c,\ p$ and $d$ to not converge due to the complex links between their values and the observational equivalence of the resulting targets.
(a) Schematic of the system.
(b) Transition parameter values fitting the observed targets well.
Figure 1: Simple example of an SIS model. Parameters $\beta$ and $\gamma$ are transition parameters. Note the density in panel (b) represents an informative prior, with arrows illustrating how it impacts upon the joint posterior and hence the marginal posterior of $\gamma$. The annotated density on the x axis for $\beta$ illustrates how the information in that parameter restricts the joint posterior, and thus provides indirect information on $\gamma$, the marginal distribution of which is now affected.
Figure 2: Resulting chains of running the Metropolis-Hastings algorithm with improper uniform priors and with informative priors.
As identified by Alarid-Escudero et al. [4], the main problem in models of this type is the lack of identifiability in the parameter space, given the information available. One solution is to obtain additional data, or an additional target for which data may be available.
The approach adopted here is to add additional information about the parameters in the form of prior information. This prior information is specified by using information from other sources (e.g. from expert knowledge or previous experience of these parameters). Note that while the term ‘prior’ may in ordinary usage be associated with information available before collecting data (i.e. targets), in the Bayesian paradigm it is the fact that information other than contained in the targets may be incorporated that is important.
In this case, an informative prior is used. We started with using uniform priors for each parameter. However, only one posterior distribution (for parameter p) was distinct from the priors. The next step was to impose normal prior for two parameter $c\sim N(9,1)\text{ and }p\sim N(0.07,0.01)$ (See rhs of Figure 2 for results). The chains were compared using the Gelman and Rubin’s convergence diagnostic[5]. The no prior chains scored a point estimate 2.58 with an upper confidence interval of 5.05 and the chains with normal priors scored a 1.05 point estimate with upped C.I. of 1.15. The posterior distributions for the parameters were very close to the imposed priors, highlighting the fact that there is no identifying information in the model itself.
In summary, in this section we have illustrated how the problem of non-identifiability can arise in even a relatively simple model, and shown how the use of additional information solves this problem. It was after initial runs that the modeller will detect that non-identifiability is occurring, and by carefully considering what information to add regarding other parameters, identifiability may be achieved. In practice, such additional information may come from, for example, monitoring a number of individual cases of the disease to better quantify the time to recovery.
5 Example: HPV model
The second model that we use in this paper to illustrate the challenges of the Bayesian MCMC approach to model calibration is an individual-based model, also called an agent-based model. An individual-based model simulates the characteristics of each individual and keeps track of each individual’s history, recording all information relevant to the disease in question.
The advantages of agent based models are that they can incorporate heterogeneity (through each agent), they can be physically realistic, and they allow for interventions that impact on individuals differently. They can commonly be used when modelling dynamic aspects of infectious diseases as described for example in [6].
The challenges associated with these models is that they take more computer time to run, they can be difficult to build, and of interest to this example they can be harder to calibrate.
The example we use here is based on the human papillomavirus (HPV) which causes cervical cancer. We follow the model developed by Kim et al. [7] which considers transitions between Healthy state, HPV infection, pre-malignant changes called cervical intra-epithelial neoplasia (CIN) 1, CIN 2 and CIN 3 (which are different levels of severity) and, finally, cancer (see Figure 3 for details).
The model allows both forward and backward transitions which leads to a complex structure. Additionally, the model’s parameters vary by strain (i.e. transition probabilities to CIN states depend on the strain of HPV infection), and therefore the number of parameters for calibration is high. We have deliberately selected this HPV model for this paper because it is likely to exhibit many of the challenges that were discussed in Section 2.1.
Figure 3: Diagram of the HPV model modified from Kim et al. [7]. The arrows represent transitions between health states (formulated in the model in terms of probabilities).
5.1 Data
We assume that all of the targets have normal distributions $f_{k}\sim N(d_{k},\sigma_{k}),$ centred at the observed data $d_{k}$ and with variances $\sigma_{k}$ that were calculated from the data. We assume all targets to be independent. Because of that assumption, our likelihood is just a product of individual likelihoods
$f(\boldsymbol{x})=\prod_{k=1}^{K}f_{k}(x_{k})=\prod_{k=1}^{K}\frac{1}{\sigma_{k}\sqrt{2\pi}}e^{-\frac{1}{2}\left(\frac{x_{k}-d_{k}}{\sigma_{k}}\right)^{2}},$
where $K=31$ is the number of calibration targets and $\boldsymbol{x}=(x_{1},\ldots,x_{K})$ are the model predictions. This product of likelihoods is scaled by each distribution probability density function at the mode and transformed onto a log scale to create a simplified goodness of fit (GOF) measure in the following way
$GOF_{total}=-2\sum_{k=1}^{K}\ln\left[\frac{f_{k}(x_{k})}{f_{k}(d_{k})}\right].$
5.2 Calibration using Metropolis-Hastings
In this section, we present some results of applying the Metropolis-Hastings algorithm to the HPV model calibration. They illustrate some of the problems associated with calibrating complex models using the Metropolis-Hastings algorithm described in Section 2.1.
The trade-off between adequate mixing and good acceptance rate is a significant source of difficulty. We have considered many different configurations of the algorithm in the process of calibrating the HPV model. We present three choices in the remainder of this section.
5.2.1 The Metropolis-Hastings algorithm with independent normal proposals
The first attempt at calibration was made where all parameters were subject to random changes on each step of the algorithm (i.e. $s=0$ case in Appendix Algorithm 1). This, however, is a less than ideal approach. As was illustrated in Figure 1(a) of Section 2, when you have two competing transition probabilities, successfully calibrating both can be impossible if the data available for calibration is population-averaged. Unfortunately, in our case, the only data available were the distributions of population-level observations. This is a common problem of incompatibility of population-level data (the most commonly collected and available) with agent-based models. As a necessary restriction, all backward transition probabilities were constrained at high-likelihood values for all subsequent realisations of the algorithm.
To further address the problem of low acceptance rate, we utilise the $s$ parameter in Appendix Algorithm 1. We will refer to this case as the ‘partial independent normal proposals’. At each step of the algorithm, we select $s$ parameters to update, while the remaining $n-s$ parameters are set to their values at the previous step. It is intuitively obvious that the algorithm is more likely to make a reasonable proposal (at random) for fewer parameters, comparing to the full parameter space of $n$ parameters. The number of parameters proposed to be updated was the same at each step. The main benefit of this approach was to increase the acceptance rate to $\sim$15%, compared to $<$1%.
5.2.2 The Metropolis-Hastings algorithm with informative priors
To further improve the performance of Appendix Algorithm 1 and particularly target the problem of the non-identifiability of the parameters, we introduce informative priors. There are many ways to incorporate any information available at a given time into priors. If the priors for each parameter are assumed to be independent, then the prior $\pi(\boldsymbol{x})$ in Appendix Algorithm 1 is just a product of one-dimensional priors $\pi(\boldsymbol{x})=\prod_{i=1}^{n}\pi_{i}(x_{i}),$ where each $\pi_{i}$ is a prior distribution for $i$’s parameter. The choice of priors represents our beliefs about parameters without taking new evidence into account. Gamma distributions were chosen to be fitted to the parameters because of their $[0,\infty)$ support and their shape. Priors on the model’s parameters were imposed to limit the algorithm’s search to (what we know to be) a plausible region of the space. The uniqueness of such a region can depend on many things. Possible sources of information from which one could construct priors also include expert opinions and alternative data sets.
Having informative priors produced the best results for the HPV model calibration. We have chosen to present only those results in detail in the next section, focusing on demonstrating the effects of having an informative prior on a complex parameter space with correlated parameters.
6 Algorithm results with informative priors
As highlighted above, the prior distribution can be thought of as ‘other information’ about the parameters than is provided by the likelihood. In this case, structural priors can be placed upon the parameter space by observing that large sections are not compatible with plausible data values. As highlighted in Section 5.2.1, plausible values had also to be chosen for the backward transitions.
To improve the convergence of chains and address the problems described in Section 2.1, structural priors on the joint parameter space were imposed. The chains with the best GOF were chosen out of all the chains obtained using all previous attempts described in Section 5.2. Those chains, containing the information on plausible parameter values, were used to fit the priors. The gamma distributions were fitted so that $x_{i}\sim\Gamma(\widehat{\alpha}_{i},\widehat{\beta}_{i})$, where $x_{i}$ is the $i^{th}$ parameter and $\widehat{\alpha}_{i}$ and $\widehat{\beta}_{i}$ were estimated from the identified chain.
Additionally, instead of updating all parameters at once, on each step we randomly chose a subset of the parameter space of cardinality $s=7$ to be updated. This number was chosen on a trial-and-error basis and was constant throughout the implementation of the algorithm. The 7 parameters to be updated were chosen randomly at each step. That increased the acceptance rate of the algorithm and improved the mixing of the chain produced by Appendix Algorithm 1.
The resulting chains of the Metropolis-Hastings algorithm with informative priors and partial independent normal proposals for four selected parameters can be found in Figure 4 along with their densities. Although here we focus on four of the most interesting parameters, the resulting densities for all 26 parameters are presented in the Appendix Figure 1. It is visible from the density plots in Figure 4 that some multi-modality exists in the space. All three chains are recognising the multi-modality and sampling from the same distribution. This multi-modality was not picked up by the original chain, chosen to inform the gamma priors. This highlights the importance of prior elicitation in complex models. The balance between informing some parameters from the data or informing them through the priors could be drastically shifted towards the priors in a situation of non-identifiability[4] or just general lack of information in the available data.
Figure 4: Results of running the Metropolis-Hastings algorithm with informative priors and partial independent normal proposals for four selected parameters. For each parameter, both the chains and distributions are presented. Each parameter represents a multiplier to a baseline monthly probability of transitioning from one state to another, conditional on a type of infection (see Figure 3 for details). The conditioning on the infection type is presented in parentheses. See Figure 3 and Kim et al. [7] for details on baseline probabilities of transitioning.
Despite the complications associated with the application of the Metropolis-Hastings algorithm to the HPV model calibration, it is noted that when compared to the original method of Kim et al. [7] the GOF was consistently better for all chains (low scores being better), staying near 20, fluctuating up and down around that value. Despite the problems, this was true for all approaches, even the ones that we have chosen not to present here. If we compare this to the simple Monte Carlo approach of Kim et al. [7], the best fitting parameter values had GOF values around 122 for the Irish data (with minimal value at 72 and maximal value at 185 in a sample of 133 best fitting values). More details can be found in the Appendix.
As was mentioned in Section 5.1, the likelihood is a product of independent normal density functions and GOF is the scaled likelihood on the log scale. The details of the means and standard deviations can be found in Appendix Table 1. If we simplify the GOF measure, we can see that it is $\chi^{2}$ with 31 degrees of freedom (assuming independence).
$\begin{split}GOF_{total}&=-2\sum_{k=1}^{31}\ln\left[\frac{f_{k}(x_{k})}{f_{k}(d_{k})}\right]\\ &=-2\sum_{k=1}^{31}\ln\left(\frac{\frac{1}{\sqrt{2\pi\sigma_{k}^{2}}}\cdot e^{-\frac{(x_{k}-d_{k})^{2}}{2\sigma_{k}^{2}}}}{\frac{1}{\sqrt{2\pi\sigma_{k}^{2}}}\cdot e^{-\frac{(d_{k}-d_{k})^{2}}{2\sigma_{k}^{2}}}}\right)\\ &=\sum_{k=1}^{31}\frac{(x_{k}-d_{k})^{2}}{\sigma_{k}^{2}}=\sum_{k=1}^{31}\left(\frac{x_{k}-d_{k}}{\sigma_{k}}\right)^{2}\sim\chi^{2}_{k=31}.\end{split}$
The corresponding p-values are 93.58% and $\ll 0.01\%$ for our approach (GOF=20) and Kim et al. [7] approach (GOF=122) respectively. However, the fit of our approach appears to be almost ‘too good’ relative to $\chi^{2}_{31}$. That can be explained by the correlation in the targets, so if one is fitting well, a few others are also fitting well. If we remove the assumption of independence, that measure of GOF would have a different distribution (depending on the correlations), but the mean of that distribution will be shifted towards zero compared to $E(\chi^{2}_{k=31})=31$. This will make the value of GOF=20 more realistic.
In general, the parameter values identified by the original approach were less concentrated, compared to posterior distributions produced by the Metropolis-Hastings algorithm. Most of the ‘best fitting’ values spanned the entire search range and didn’t identify areas with higher likelihood. Additionally, some of the good-fitting values identified by the Metropolis-Hastings algorithm were outside of the originally specified ranges which was also missed by the original approach. Overall, the Metropolis-Hastings algorithm was able to find better-fitting parameter combinations, the distributions identified had smaller variances and the search required fewer model runs to produce those results.
7 Discussion
Given the advantages of a Bayesian stochastic perspective, and in particular here an MCMC approach, the fact that formal convergence is not always easily obtained without substantial additional work requires much attention. Such work includes better data collection or data representation in terms of priors as well as tuning the MCMC to the problem.
A key message of this work is that good prior elicitation in complex models is essential for a good fit of the model’s parameters. While here we have examined as an outcome the improvement in GOF, this must be considered in light of the expertise required to apply these methods to model calibration. Care is needed with starting values and proposal distributions to ensure that the chains have converged and that the parameter space is properly explored.
The Bayesian approach to model calibration is becoming more popular[8] and more research is focusing on MCMC techniques in particular. Several calibration approaches become more difficult to apply in the context of stochastic models, such as stochastic microsimulations or agent-based models[8].
One of the main observations in this work was that it is possible to explicitly incorporate additional information into the inference in the form of priors during the calibration process and that this had a significant effect on the algorithm’s performance. This has a big impact on the issue of non-identifiability. Once the additional information was included (here through the revision of the prior), the algorithm was not impaired by the presence of some local multi-modality and all three chains sampled from the same distribution. This led to a desirable effect of convergence that was coupled with the improved acceptance rate of the algorithm and visible lack of autocorrelation in the sample.
The fact that the application of the Metropolis-Hastings algorithm yielded results with a lower GOF than the simple Monte Carlo approach is not surprising, given that the Metropolis-Hastings algorithm is designed to lead the sampler to a high likelihood area, whether it is a global or a local maximum. Additionally, this technique provides probabilistic estimates of the parameters of interest and reflects their joint uncertainty. Usually, the acceptance criterion is applied after the fact, which leads to inefficiency.
In a population-level model of Hepatitis C infection a similar process of calibration was performed [9]. This case had none of the problems that we had with the HPV model calibration. This was mostly because good priors were provided by an expert. Furthermore, because the Hepatitis C model did not have any backward transitions, there were fewer problems with identifiability, compared to the HPV model. Another possible advantage of that model was its relative simplicity, making it difficult to over-parameterise, compared to the individual-based models. This allowed us to try multiple strategies, before selecting the best one. We hypothesise that the relatively simpler structure of the Hepatitis C model, compared with the HPV model, yields a smoother probabilistic link between the model parameters and outputs. Of course, model choice is very influential in model calibration and different model types might yield different calibration results.
Funding
This work was supported by the Science Foundation Ireland [grant number 12/IA/1683] and the Health Research Board [grant number RL13/04].
Acknowledgements
We would like to thank MACSI (University of Limerick) and the Health Research Board who have provided the funding for this research. We are also grateful to Jane Kim for sharing the code from their paper[7]. Daria Semochkina would also like to thank Paul Smith for his advice on article preparation and many helpful discussions.
Appendix A Appendices
A.1 Metropolis-Hastings algorithm
The central idea of the Metro[polius-Hastings algorithm is to update locally, i.e. use the accepted parameter value $\boldsymbol{x}_{(t)}$ at time step $t$ to generate the next proposed value. This makes it easier to think of a suitable proposal step (that is, of course, conditional), albeit with the disadvantage of producing a Markov chain as opposed to an independent realisation. This Markov chain converges to the posterior distribution of interest after enough steps. The sample values are taken from the Markov chain after convergence. Below we present the M-H algorithm we used to calibrate the HPV model.
Algorithm 1 HPV model calibration using Metropolis-Hastings.
Starting from $\boldsymbol{x}^{(0)}=(x^{(0)}_{1},x^{(0)}_{2},...,x^{(0)}_{n})$ for $t=1$ to $N=N_{steps}$ do
1:Draw $\boldsymbol{x}\sim q(\cdot|\boldsymbol{x}^{(t-1)}),$ where $q$ is the proposal distribution
2:For some $s<n$ randomly select $n-s$ parameters $x_{j_{1}},...x_{j_{n-s}}$ and set them to previous values $x_{j_{1}}=x^{(t-1)}_{j_{1}},...,x_{j_{n-s}}=x^{(t-1)}_{j_{n-s}}$
3:Run the model for a cohort of 25,000 women, given the new parameter set $\boldsymbol{x}$
4:Calculate the likelihood of the newly proposed step $f(\boldsymbol{x})$
5:Compute
$p(\boldsymbol{x}|\boldsymbol{x}^{(t-1)})=\min\left\{1,\frac{\pi(\boldsymbol{x})f(\boldsymbol{x})q(\boldsymbol{x}^{(t-1)}|\boldsymbol{x})}{\pi(\boldsymbol{x}^{(t-1)})f(\boldsymbol{x}^{(t-1)})q(\boldsymbol{x}|\boldsymbol{x}^{(t-1)})}\right\},$ (1)
where $\pi(\boldsymbol{x})$ is the prior distribution
6:With probability $p(\boldsymbol{x}|\boldsymbol{x}^{(t-1)})$ accept the step and set $\boldsymbol{x}^{(t)}=\boldsymbol{x}$, otherwise reject the step and let $\boldsymbol{x}^{(t)}=\boldsymbol{x}^{(t-1)}$
End
Here $q$ is the proposal distribution, which is often difficult to choose. The probability of acceptance in (1) does not depend on the normalising constants for either the proposal distribution $q$ or the distribution of interest $f$.
A.2 HPV model
In this appendix, we present further details about the HPV model, described in Section 5. In Appendix Table 2 we present the best five parameter combination that produced the lowest GOF out of all the chains for the two different proposal distribution strategies described in Section 5.2.1 and Section 5.2.2. We chose five combinations, instead of presenting mean and variances, to demonstrate that there is still a non-identifiability issue present.
Calibration Target $d_{k}$ $(\sigma_{k})$
Duration of HPV infection
 Low risk, $<$30 years 9.775 (0.599)
 High risk (other), $<$30 years 11.720 (0.638)
 High risk 16, $<$30 years 12.905 (1.334)
 High risk 18, $<$30 years 10.015 (1.620)
 Low risk, $>$30 years 9.910 (0.378)
 High risk (other), $>$30 years 11.300 (0.622)
 High risk 16, $>$30 years 10.875 (1.099)
 High risk 18, $>$30 years 10.960 (1.857)
Prevalence of high-risk HPV
  infection
 20-24 years 0.398 (0.107)
 25-34 years 0.224 (0.069)
 35-44 years 0.108 (0.044)
 45-54 years 0.077 (0.036)
 55-64 years 0.060 (0.027)
Distribution of HPV types
  among women with CIN1
 High risk 16 0.2222 (0.025)
 High risk (other) 0.1025 (0.022)
Distribution of HPV types
  among women with CIN23
 High risk 16 0.6333 (0.148)
 High risk 18 0.0761 (0.010)
 High risk (other) 0.0557 (0.055)
Distribution of HPV types
  among women with cancer
 High risk 16 0.6810 (0.029)
 High risk 18 0.1067 (0.010)
Incidence rate of invasive cancer
   (cases per 100,000 per year)
 25-29 years 6.34 (2.67)
 30-34 years 14.67 (3.60)
 35-39 years 20.13 (4.58)
 40-44 years 23.83 (3.74)
 45-49 years 21.83 (4.28)
 50-54 years 18.29 (5.06)
 55-59 years 18.25 (4.34)
 60-64 years 13.47 (3.83)
 65-69 years 15.59 (4.08)
 70-74 years 14.46 (4.07)
 75-79 years 16.38 (6.74)
Appendix Table 1: Calibration targets with corresponding distribution parameters ($d_{k}$ is the mean and $\sigma_{k}$ is the standard deviation). The mean and variance for each parameter were elicited from the available data independently for lack of a better alternative. For full details see section 5.1.
We opted not to use the calibration ranges adopted by Kim et al. [7]. The Metropolis-Hastings algorithm leads the search in the direction of good-fitting values. Consequently, we do not need to worry about our computational budget to be able to cover the entire space. It is apparent that although some of the parameters are comparable with the results in Kim et al. [7], there is variation even between the two approaches presented in this paper. Transitions from HPV to CIN1 (HR 16); CIN1 to CIN23 (LR); CIN23 to CA (HR other) and CIN23 to CA (HR 18) have much higher values with the independent proposals strategy, compared to the strategy with gamma priors. Immune Degree (LR) has very different values in the two different strategies (this was assumed to be zero in the original paper). These results illustrate some of the concepts presented in Section 2, highlighting that different parameter combinations can give similar results.
Additionally, we present posterior distributions of all parameter sets with the corresponding GOF below 25 (see Appendix Figure 1).
Partial independent normal proposals Informative priors
GOF 14.54 14.76 15.11 15.34 15.36 17.14 17.17 17.27 17.39 17.62
Immune Factor (HR other) 0.85 0.99 0.97 0.95 0.86 0.97 0.96 0.97 0.98 0.97
Immune Factor (HR 16) 0.94 0.97 0.98 0.94 0.95 0.97 0.98 0.98 0.97 0.97
Immune Factor (HR 18) 0.29 0.13 0.07 0.12 0.27 0.89 0.88 0.88 0.89 0.88
Immune Degree (LR) 0.77 0.45 0.96 0.90 0.75 0.12 0.10 0.11 0.11 0.09
Immune Degree (HR other) 0.31 0.60 0.64 0.93 0.48 0.12 0.17 0.18 0.16 0.23
Immune Degree (HR 16) 0.18 0.03 0.16 0.03 0.15 0.90 0.88 0.89 0.89 0.90
Immune Degree (HR 18) 0.24 0.19 0.42 0.13 0.25 0.76 0.73 0.75 0.75 0.76
Healthy to HPV (LR) 0.77 0.51 2.94 0.66 0.77 4.86 5.09 4.67 4.43 4.97
Healthy to HPV (HR other) 0.60 3.48 0.97 0.31 0.55 5.41 5.93 5.55 5.46 5.60
Healthy to HPV (HR 16) 0.08 0.10 0.05 0.06 0.08 6.71 6.15 6.12 6.03 6.73
Healthy to HPV (HR 18) 5.17 3.84 5.18 5.20 5.17 2.98 2.60 2.60 2.90 3.02
HPV to CIN1 (LR) 6.22 8.53 0.59 2.86 6.18 3.43 3.27 3.47 3.73 3.37
HPV to CIN1 (HR other) 0.44 0.25 0.44 1.16 0.45 0.22 0.25 0.17 0.22 0.25
HPV to CIN1 (HR 16) 22.98 21.33 33.48 25.80 23.35 3.63 3.65 3.65 3.65 3.65
HPV to CIN1 (HR 18) 0.06 0.26 0.15 0.12 0.06 0.18 0.14 0.09 0.16 0.19
HPV to CIN23 (LR) 0.21 0.18 0.10 0.09 0.21 0.04 0.03 0.04 0.04 0.04
HPV to CIN23 (HR other) 0.04 0.03 0.09 0.08 0.05 0.02 0.02 0.02 0.02 0.02
HPV to CIN23 (HR 16) 0.88 0.56 0.89 0.10 0.84 0.79 0.80 0.78 0.80 0.80
HPV to CIN23 (HR 18) 0.01 0.01 0.01 0.00 0.01 0.06 0.07 0.06 0.06 0.06
CIN1 to CIN23 (LR) 12.13 20.40 31.51 24.49 12.25 3.99 4.91 4.05 4.49 4.86
CIN1 to CIN23 (HR other) 0.51 0.13 0.19 0.49 0.56 0.05 0.03 0.04 0.06 0.04
CIN1 to CIN23 (HR 16) 8.99 13.08 13.36 14.02 9.43 0.16 0.17 0.28 0.14 0.17
CIN1 to CIN23 (HR 18) 3.12 3.08 2.14 3.54 2.83 5.60 5.91 5.83 5.70 5.49
CIN23 to Local Cancer (HR other) 17.81 10.84 16.15 14.59 18.14 4.27 4.26 4.33 4.22 4.24
CIN23 to Local Cancer (HR 16) 6.04 2.52 6.91 5.50 6.06 5.02 5.22 5.16 5.14 5.10
CIN23 to Local Cancer (HR 18) 10.23 2.60 24.75 35.58 11.10 5.27 5.17 5.43 5.60 5.21
Fixed parameters
CIN23 to CIN1 (LR) 1.40 CIN23 to Healthy (LR) 2.42 CIN1 to Healthy (LR) 4.22
CIN23 to CIN1 (HR other) 3.31 CIN23 to Healthy (HR other) 4.23 CIN1 to Healthy (HR other) 0.56
CIN23 to CIN1 (HR 16) 1.38 CIN23 to Healthy (HR 16) 4.00 CIN1 to Healthy (HR 16) 4.80
CIN23 to CIN1 (HR 18) 5.63 CIN23 to Healthy (HR 18) 5.28 CIN1 to Healthy (HR 18) 12.87
CIN23 to HPV (LR) 4.38 CIN1 to HPV (LR) 2.99 HPV to Healthy (LR) 7.09
CIN23 to HPV (HR other) 2.09 CIN1 to HPV (HR other) 5.03 HPV to Healthy (HR other) 4.42
CIN23 to HPV (HR 16) 5.12 CIN1 to HPV (HR 16) 5.69 HPV to Healthy (HR 16) 7.40
CIN23 to HPV (HR 18) 1.98 CIN1 to HPV (HR 18) 4.15 HPV to Healthy (HR 18) 5.44
Appendix Table 2: Five best-fitting values for two proposal strategies: Partial independent normal proposals and Informative priors (Partial independent normal with gamma priors), described in Section 5.2.1 and Section 5.2.2. For simplicity, everywhere in this table, the parentheses imply a transition between states is conditional on the type of HPV infection specified in the parentheses. Corresponding GOF measures are also presented. See Figure 3 for details.
Appendix Figure 1: Posterior densities of all parameters with partial independent proposals, gamma priors and GOF25absent25$\leq$ 25≤ 25. See Section 5.2.2 for full details of the approach .